10r2+20r-88=-8
Could not figure out how to factor this radical out. Is it possible to have no solution?

Answer by ankor@dixie-net.com(22740)

10r2+20r-88=-8
:
10r^2 + 20r - 88 + 8 = 0
10r^2 + 20r - 80 = 0
simplify, divide both sides by 10
r^2 + 2r - 8 = 0
Factors easily to
(r+4)(r-2) = 0
Two solutions
r = -4
r = 2
:
:
Check solutions in original equation
r= -4
10(-4^2) + 20(-4) - 88 = -8
10(16) - 80 - 88 = -8
:
You can check the r=2 solution